OpenStudy (anonymous):

Solve for4x+5y=20 for y

OpenStudy (cj49):

y=20-4x/5

OpenStudy (cj49):

x=5,y=0

OpenStudy (texaschic101):

4x + 5y = 20 --- subtract 4x from both sides 5y = -4x + 20 --- divide both sides by 5 y = -4/5x + 4